From the top of a tower 100m in height a ball is dropped and at the same time another ball is projected vertically upward from the ground with velocity of 25m/s.Find when amd where will  the two balls will meet.Take g=9.8m/s2 

Grade:11

2 Answers

Himanshu Rathour
5 years ago
L
Himanshu Rathour
5 years ago
Let the two balls meet after time t andaat distance x from ground 
For ball 1
x=25t-1/2gt²
For ball 2
100-x=0t+1/2gt²
100-(25t-1/2gt²)=1/2gt²
100-25t=0
t=4s
X=25*4-1/2(9.8)(4)
x=80.4m